# Scalable Graph Generative Modeling via Substructure Sequences

### v1

@@ 3.3 Tokenization
Substructure Encoder. To embed substructure sequences (e.g., random walks), we adopt a Transformer encoder, which models global dependencies among all nodes in the sequence. As shown in Table 1, Transformer outperforms alternatives such as mean pooling, GRU [10], and GIN [68]. We attribute this to several factors: (1) mean pooling is a restricted, less expressive form of attention in Transformer; (2) GRUs suffer from optimization instability due to vanishing or exploding gradients [80]; and (3) using GIN requires converting walks into subgraphs, reintroducing message-passing and its known limitations.
Positional Embedding. We also explore adding positional embeddings (PEs) to node features—using Laplacian PE on ogbn-arxiv and random walk PE on ogbg-HIV [42]. However, we observe no consistent performance gain. This may be due to the tokenization scheme: since each token represents an entire substructure, global node positions are less relevant. Moreover, computing PEs incurs cubic time complexity in the node number [29], making them impractical for large-scale graphs.
@@ 3.5 Masking
Masking Ratio. Figure 5 shows that a high masking ratio consistently improves performance, echoing trends observed in vision [19]. This is likely because random walks often produce redundant or noisy substructures. High masking reduces such redundancy, resulting in a more challenging and informative learning signal that encourages the model to capture high-level semantics.
@@ 4 Comparisons to State-of-The-Art Models
Supervised GAT [55] 83.1 ± 0.3 91.9 ± 0.5 87.9 ± 0.5 76.9 ± 0.8 50.7 ± 0.3 72.10 ± 0.13 79.45 ± 0.59 5.7
GPM [63] 84.7 ± 0.1 92.7 ± 0.3 90.0 ± 0.4 80.2 ± 0.4 52.2 ± 0.2 72.89 ± 0.68 82.62 ± 0.39 1.3
Contrastive GCA [79] 83.3 ± 0.5 92.4 ± 0.2 87.1 ± 0.2 77.4 ± 0.1 49.0 ± 0.1 71.23 ± 0.09 78.39 ± 0.03 6.9
BGRL [48] 83.9 ± 0.3 92.5 ± 0.2 88.2 ± 0.2 77.5 ± 0.8 49.7 ± 0.2 70.51 ± 0.03 78.59 ± 0.02 5.7
CCA-SSG [73] 81.8 ± 0.5 91.8 ± 0.6 88.6 ± 0.3 75.3 ± 0.8 47.5 ± 0.2 71.24 ± 0.20 75.27 ± 0.05 8.6
Generative GraphMAE [21] 81.0 ± 0.5 92.0 ± 0.3 89.2 ± 0.5 77.1 ± 0.5 50.5 ± 0.1 71.75 ± 0.17 78.89 ± 0.01 6.0
GraphMAE 2 [22] 81.3 ± 0.4 92.4 ± 0.2 88.3 ± 0.9 77.6 ± 0.4 50.4 ± 0.1 71.89 ± 0.03 79.33 ± 0.01 5.4
S2GAE [47] 80.1 ± 0.5 91.4 ± 0.1 85.3 ± 0.1 75.3 ± 0.8 48.1 ± 0.8 67.77 ± 0.36 76.70 ± 0.03 10.3
Bandana [77] 83.5 ± 0.5 91.4 ± 0.7 87.7 ± 0.2 77.3 ± 0.3 47.9 ± 0.6 71.09 ± 0.24 77.68 ± 0.05 8.1
@@
Node Classification on homophily graphs. We evaluate G2PM on a suite of homophily graphs of varying scales, including Pubmed, Photo, Computers, WikiCS, Flickr, ogbn-arxiv, and ogbn-products (see Table 2 for dataset statistics). Pre-training is conducted on the full graph. We adopt a linear probe setup: node embeddings are frozen after pre-training and used to train a separate classifier, where we take 10/10/80 random split for Pubmed, Photo, and Computers, and the official split for the remaining datasets. Accuracy is used as the evaluation metric. We compare G2PM to supervised GAT [55], non-message-passing GPM [63], contrastive methods (GCA [79], BGRL [48], CCA-SSG [73]), and generative methods (GraphMAE [21], GraphMAE 2 [22], S2GAE [47], Bandana [77]).
@@
Supervised GIN [68] 75.8 ± 0.8 70.3 ± 0.3 57.7 ± 0.8 74.4 ± 0.9 83.4 ± 0.6 73.3 ± 0.5 39.4 ± 1.4 6.3
GPM [63] 77.0 ± 0.9 75.1 ± 0.3 59.0 ± 0.0 74.6 ± 1.4 82.4 ± 0.3 82.7 ± 0.5 43.1 ± 0.3 3.0
Contrastive GraphCL [70] 75.5 ± 0.3 72.4 ± 2.1 57.3 ± 0.9 68.3 ± 2.6 82.9 ± 0.3 71.1 ± 0.4 37.9 ± 2.4 8.0
JOAO [71] 76.8 ± 0.3 73.4 ± 1.5 58.5 ± 0.5 72.3 ± 1.0 82.2 ± 0.3 70.2 ± 3.1 39.9 ± 0.6 6.0
MVGRL [18] 75.7 ± 0.7 70.4 ± 2.1 60.5 ± 0.6 71.5 ± 1.2 83.6 ± 0.2 74.2 ± 0.7 39.5 ± 1.8 5.7
InfoGCL [67] 77.3 ± 0.6 74.6 ± 0.7 58.7 ± 0.7 73.4 ± 1.0 80.3 ± 0.7 75.1 ± 0.9 39.3 ± 0.5 5.4
Generative GraphMAE [21] 77.8 ± 0.9 73.2 ± 1.4 60.6 ± 0.0 73.7 ± 0.8 84.8 ± 0.5 75.5 ± 0.7 37.6 ± 2.5 4.1
S2GAE [47] 75.6 ± 0.8 72.9 ± 0.0 58.0 ± 0.9 71.6 ± 0.8 80.6 ± 0.4 75.8 ± 0.6 37.9 ± 1.8 7.0
@@
Graph Classification. We evaluate G2PM on seven datasets: five molecular graphs (HIV, PCBA, SIDER, MUV, ClinTox) and two social networks (IMDB-B, Reddit-M12K). Dataset statistics and results are summarized in Table 3. We use public splits for molecular graphs and 80/10/10 random splits for social networks, following a linear probe protocol. Baselines include supervised GAT [55], non-message-passing GPM [63], contrastive methods (GraphCL [70], JOAO [71], MVGRL [18], InfoGCL [67]), and generative models (GraphMAE [21], S2GAE [47]). G2PM consistently achieves the best performance across all datasets. In contrast, the model without pre-training performs worst on average, highlighting the importance of capturing domain-specific substructure distributions to support discriminative generalization.
@@ 5 Cross-Domain Graph Learning
Cross-Domain Transfer. We evaluate the cross-domain transferability of G2PM in Table 4, measuring generalization under distribution shifts across domains and tasks. In this setting, the model is pre-trained on a source graph and fully fine-tuned on a target graph. Since feature spaces differ across graphs, we introduce a learnable linear projection layer before the pre-trained encoder, which is jointly finetuned during transfer. Baselines include GNNs (GAT [55] for ogbn-arxiv and GIN [68] for HIV), as well as GPM [63], BGRL [48], and GraphMAE [21]. G2PM achieves the best transfer results, showing positive gains in 3 out of 4 setups. In contrast, message-passing methods only transfer well across closely related domains (e.g., HIV $\to$ PCBA). We attribute this to the ability to learn transferable structural patterns, enabling generalization even across domain and task boundaries (e.g., Arxiv $\to$ HIV), while message-passing remains sensitive to subtle structural shifts [61].
@@
GNN [55, 68] 78.3 (1.2 $\downarrow$) 70.1 (5.7 $\downarrow$) 71.1 (1.0 $\downarrow$) 71.9 (1.6 $\uparrow$)
GPM [63] 82.0 (0.6 $\downarrow$) 74.3 (2.7 $\downarrow$) 71.4 (1.5 $\downarrow$) 76.4 (1.3 $\uparrow$)
BGRL [48] 78.8 (0.2 $\uparrow$) 72.5 (3.8 $\downarrow$) 68.6 (1.9 $\downarrow$) 72.9 (0.6 $\downarrow$)
GraphMAE [21] 77.5 (1.4 $\downarrow$) 74.7 (3.1 $\downarrow$) 69.9 (1.9 $\downarrow$) 73.4 (0.2 $\uparrow$)
@@
Table 5: Cross-domain pre-training results on text-attributed graphs processed by [32], where node features are aligned via a textual encoder.
@@
BGRL [48] 70.8 ± 0.2 86.5 ± 0.3 68.5 ± 1.6
GraphMAE [21] 70.3 ± 0.3 87.8 ± 0.4 64.1 ± 0.5
OFA [32] 71.4 ± 0.3 84.7 ± 1.3 72.0 ± 1.6
GFT [60] 71.9 ± 0.1 89.3 ± 0.2 72.3 ± 2.0
@@
Cross-Domain Pre-Training. Table 5 presents results on cross-domain pre-training using three diverse graph types: Arxiv (academic network), FB15K237 (knowledge graph), and HIV (molecular graph) [32]. All graphs are text-attributed; we use a shared textual encoder to project node descriptions into a unified embedding space, enabling a single model to operate across domains. We compare G2PM to pre-training baselines (BGRL [48], GraphMAE [21]) and graph foundation models (OFA [32], GFT [60]). G2PM achieves the best or second-best performance across all datasets, highlighting its superior ability to learn transferable substructure representations, even under significant domain shifts—where message-passing-based methods often struggle.
@@ 6 Related Works
Graph Transformers. Inspired by the success of Transformers in vision and language [11, 7, 52, 19, 12], several works have extended this architecture to graphs [29, 13, 42, 20, 9, 8, 74, 75]. Typically, graph Transformers treat nodes as tokens and apply self-attention over all pairs. However, the quadratic complexity of node attention limits their scalability, making them impractical for large graphs [65]. To address this, recent works propose tokenizing graphs via substructures such as random walks or motifs [69, 23], enabling sequence-based modeling [63, 9, 20]. For example, GPM [63] applies a ViT-style architecture over substructure sequences. Building on this idea, our work explores generative pretraining using a Transformer over substructures, entirely removing message passing.
Generative Pretraining. Generative pretraining has driven major advances in vision and language. These methods predict masked content from visible context, enabling learning from large-scale unlabeled data and powering modern foundation models [52]. This paradigm has extended to other domains (e.g., video, audio, biology), but its impact on graphs remains limited. More recent works apply generative approaches to graphs (e.g., VGAE [27], GraphMAE [21], GraphMAE2 [22], MaskGAE [30]), but focus on reconstructing low-level signals such as nodes or links, and typically rely on message-passing GNNs, limiting their expressiveness and scalability. In contrast, G2PM introduces a fully Transformer-based, message-passing-free framework that models high-level semantic substructures, unlocking better scalability and generalization.
Random Walks on Graphs. Random walks have been widely used to represent substructures, from early unsupervised embeddings [39, 16, 69, 23] to recent deep models [63, 76, 24, 36, 50]. These methods have been shown to break key limitations of message passing, improving expressiveness [64, 37], modeling long-range dependencies [36, 50], and mitigating over-smoothing and over-squashing [57]. However, these random walk-based methods often focus on global patterns while overlooking localized structures [50]. G2PM uses random walks to sample substructures and design the MSM task to automatically balance the contributions between localized and globalized substructures.

### v3

@@ 3.3 Tokenization
Substructure Encoder. To embed substructure sequences (e.g., random walks), we adopt a Transformer encoder, which models global dependencies among all nodes in the sequence. As shown in Table 1, Transformer outperforms alternatives such as mean pooling, GRU [chung2014empirical], and GIN [xu2018how]. We attribute this to several factors: (1) mean pooling is a restricted, less expressive form of attention in Transformer; (2) GRUs suffer from optimization instability due to vanishing or exploding gradients [zucchet2024recurrent]; and (3) using GIN requires converting walks into subgraphs, reintroducing message-passing and its known limitations.
Positional Embedding. We also explore adding positional embeddings (PEs) to node features—using Laplacian PE on ogbn-arxiv and random walk PE on ogbg-HIV [rampasek2022recipe]. However, we observe no consistent performance gain. This may be due to the tokenization scheme: since each token represents an entire substructure, global node positions are less relevant. Moreover, computing PEs incurs cubic time complexity in the node number [kreuzer2021rethinking], making them impractical for large-scale graphs.
@@ 3.5 Masking
Masking Ratio. Figure 5 shows that a high masking ratio consistently improves performance, echoing trends observed in vision [he2022masked]. This is likely because random walks often produce redundant or noisy substructures. High masking reduces such redundancy, resulting in a more challenging and informative learning signal that encourages the model to capture high-level semantics.
@@ 4 Comparisons to State-of-The-Art Models
Supervised GAT [velickovic2018graph] 83.1 ± 0.3 91.9 ± 0.5 87.9 ± 0.5 76.9 ± 0.8 50.7 ± 0.3 72.10 ± 0.13 79.45 ± 0.59 5.7
GPM [wang2025neural] 84.7 ± 0.1 92.7 ± 0.3 90.0 ± 0.4 80.2 ± 0.4 52.2 ± 0.2 72.89 ± 0.68 82.62 ± 0.39 1.3
Contrastive GCA [zhu2021graph] 83.3 ± 0.5 92.4 ± 0.2 87.1 ± 0.2 77.4 ± 0.1 49.0 ± 0.1 71.23 ± 0.09 78.39 ± 0.03 6.9
BGRL [thakoor2022largescale] 83.9 ± 0.3 92.5 ± 0.2 88.2 ± 0.2 77.5 ± 0.8 49.7 ± 0.2 70.51 ± 0.03 78.59 ± 0.02 5.7
CCA-SSG [zhang2021canonical] 81.8 ± 0.5 91.8 ± 0.6 88.6 ± 0.3 75.3 ± 0.8 47.5 ± 0.2 71.24 ± 0.20 75.27 ± 0.05 8.6
Generative GraphMAE [hou2022graphmae] 81.0 ± 0.5 92.0 ± 0.3 89.2 ± 0.5 77.1 ± 0.5 50.5 ± 0.1 71.75 ± 0.17 78.89 ± 0.01 6.0
GraphMAE 2 [hou2023graphmae2] 81.3 ± 0.4 92.4 ± 0.2 88.3 ± 0.9 77.6 ± 0.4 50.4 ± 0.1 71.89 ± 0.03 79.33 ± 0.01 5.4
S2GAE [tan2023s2gae] 80.1 ± 0.5 91.4 ± 0.1 85.3 ± 0.1 75.3 ± 0.8 48.1 ± 0.8 67.77 ± 0.36 76.70 ± 0.03 10.3
Bandana [zhao2024masked] 83.5 ± 0.5 91.4 ± 0.7 87.7 ± 0.2 77.3 ± 0.3 47.9 ± 0.6 71.09 ± 0.24 77.68 ± 0.05 8.1
@@
Node Classification on homophily graphs. We evaluate G2PM on a suite of homophily graphs of varying scales, including Pubmed, Photo, Computers, WikiCS, Flickr, ogbn-arxiv, and ogbn-products (see Table 2 for dataset statistics). Pre-training is conducted on the full graph. We adopt a linear probe setup: node embeddings are frozen after pre-training and used to train a separate classifier, where we take 10/10/80 random split for Pubmed, Photo, and Computers, and the official split for the remaining datasets. Accuracy is used as the evaluation metric. We compare G2PM to supervised GAT [velickovic2018graph], non-message-passing GPM [wang2025neural], contrastive methods (GCA [zhu2021graph], BGRL [thakoor2022largescale], CCA-SSG [zhang2021canonical]), and generative methods (GraphMAE [hou2022graphmae], GraphMAE 2 [hou2023graphmae2], S2GAE [tan2023s2gae], Bandana [zhao2024masked]).
@@
Table 3: Link prediction results on three widely-used benchmarks.
Cora Pubmed ogbl-Collab
# Nodes 2,708 19,717 235,868
# Edges 10,556 88,648 2,570,930
Metric Hit@20 Hit@20 Hit@50
GCN [kipf2017semisupervised] 84.1 ± 1.2 85.1 ± 3.8 44.8 ± 1.1
GraphMAE [hou2022graphmae] 87.6 ± 1.4 87.1 ± 3.1 46.5 ± 0.9
G2PM 90.4 ± 0.8 88.0 ± 3.8 47.1 ± 0.6
Link Prediction. To further demonstrate the effectiveness of our pretraining approach, we follow wang2025neural to show the link prediction results. We use three datasets, including Cora, Pubmed, and ogbl-collab, and follow a standard 80/5/15 split for training, validation, and test sets. For evaluation, we report Hit@20 on Cora and Pubmed, and Hit@50 on ogbl-collab. The results are summarized in Table 3, where we compare our method G2PM with a basic supervised GCN [kipf2017semisupervised] and a widely used graph pretraining method, GraphMAE [hou2022graphmae]. As shown, GraphMAE significantly outperforms GCN, benefiting from its ability to utilize unlabeled data during pretraining. Notably, G2PM achieves the best performance across all datasets, likely due to its capacity to model substructures that capture latent connectivity patterns between nodes.
@@
Supervised GIN [xu2018how] 75.8 ± 0.8 70.3 ± 0.3 57.7 ± 0.8 74.4 ± 0.9 83.4 ± 0.6 73.3 ± 0.5 39.4 ± 1.4 6.3
GPM [wang2025neural] 77.0 ± 0.9 75.1 ± 0.3 59.0 ± 0.0 74.6 ± 1.4 82.4 ± 0.3 82.7 ± 0.5 43.1 ± 0.3 3.0
Contrastive GraphCL [you2020graph] 75.5 ± 0.3 72.4 ± 2.1 57.3 ± 0.9 68.3 ± 2.6 82.9 ± 0.3 71.1 ± 0.4 37.9 ± 2.4 8.0
JOAO [you2021graph] 76.8 ± 0.3 73.4 ± 1.5 58.5 ± 0.5 72.3 ± 1.0 82.2 ± 0.3 70.2 ± 3.1 39.9 ± 0.6 6.0
MVGRL [hassani2020contrastive] 75.7 ± 0.7 70.4 ± 2.1 60.5 ± 0.6 71.5 ± 1.2 83.6 ± 0.2 74.2 ± 0.7 39.5 ± 1.8 5.7
InfoGCL [xu2021infogcl] 77.3 ± 0.6 74.6 ± 0.7 58.7 ± 0.7 73.4 ± 1.0 80.3 ± 0.7 75.1 ± 0.9 39.3 ± 0.5 5.4
Generative GraphMAE [hou2022graphmae] 77.8 ± 0.9 73.2 ± 1.4 60.6 ± 0.0 73.7 ± 0.8 84.8 ± 0.5 75.5 ± 0.7 37.6 ± 2.5 4.1
S2GAE [tan2023s2gae] 75.6 ± 0.8 72.9 ± 0.0 58.0 ± 0.9 71.6 ± 0.8 80.6 ± 0.4 75.8 ± 0.6 37.9 ± 1.8 7.0
@@
Graph Classification. We evaluate G2PM on seven datasets: five molecular graphs (HIV, PCBA, SIDER, MUV, ClinTox) and two social networks (IMDB-B, Reddit-M12K). Dataset statistics and results are summarized in Table 4. We use public splits for molecular graphs and 80/10/10 random splits for social networks, following a linear probe protocol. Baselines include supervised GAT [velickovic2018graph], non-message-passing GPM [wang2025neural], contrastive methods (GraphCL [you2020graph], JOAO [you2021graph], MVGRL [hassani2020contrastive], InfoGCL [xu2021infogcl]), and generative models (GraphMAE [hou2022graphmae], S2GAE [tan2023s2gae]). G2PM consistently achieves the best performance across all datasets. In contrast, the model without pre-training performs worst on average, highlighting the importance of capturing domain-specific substructure distributions to support discriminative generalization.
@@ 5 Cross-Domain Graph Learning
Cross-Domain Transfer. We evaluate the cross-domain transferability of G2PM in Table 5, measuring generalization under distribution shifts across domains and tasks. In this setting, the model is pre-trained on a source graph and fully fine-tuned on a target graph. Since feature spaces differ across graphs, we introduce a learnable linear projection layer before the pre-trained encoder, which is jointly finetuned during transfer. Baselines include GNNs (GAT [velickovic2018graph] for ogbn-arxiv and GIN [xu2018how] for HIV), as well as GPM [wang2025neural], BGRL [thakoor2022largescale], and GraphMAE [hou2022graphmae]. G2PM achieves the best transfer results, showing positive gains in 3 out of 4 setups. In contrast, message-passing methods only transfer well across closely related domains (e.g., HIV $\to$ PCBA). We attribute this to the ability to learn transferable structural patterns, enabling generalization even across domain and task boundaries (e.g., Arxiv $\to$ HIV), while message-passing remains sensitive to subtle structural shifts [wang2024tackling].
@@
GNN [velickovic2018graph, xu2018how] 78.3 (1.2 $\downarrow$) 70.1 (5.7 $\downarrow$) 71.1 (1.0 $\downarrow$) 71.9 (1.6 $\uparrow$)
GPM [wang2025neural] 82.0 (0.6 $\downarrow$) 74.3 (2.7 $\downarrow$) 71.4 (1.5 $\downarrow$) 76.4 (1.3 $\uparrow$)
BGRL [thakoor2022largescale] 78.8 (0.2 $\uparrow$) 72.5 (3.8 $\downarrow$) 68.6 (1.9 $\downarrow$) 72.9 (0.6 $\downarrow$)
GraphMAE [hou2022graphmae] 77.5 (1.4 $\downarrow$) 74.7 (3.1 $\downarrow$) 69.9 (1.9 $\downarrow$) 73.4 (0.2 $\uparrow$)
@@
Table 6: Cross-domain pre-training results on text-attributed graphs processed by [liu2024one], where node features are aligned via a textual encoder.
@@
BGRL [thakoor2022largescale] 70.8 ± 0.2 86.5 ± 0.3 68.5 ± 1.6
GraphMAE [hou2022graphmae] 70.3 ± 0.3 87.8 ± 0.4 64.1 ± 0.5
OFA [liu2024one] 71.4 ± 0.3 84.7 ± 1.3 72.0 ± 1.6
GFT [wang2024gft] 71.9 ± 0.1 89.3 ± 0.2 72.3 ± 2.0
@@
Cross-Domain Pre-Training. Table 6 presents results on cross-domain pre-training using three diverse graph types: Arxiv (academic network), FB15K237 (knowledge graph), and HIV (molecular graph) [liu2024one]. All graphs are text-attributed; we use a shared textual encoder to project node descriptions into a unified embedding space, enabling a single model to operate across domains. We compare G2PM to pre-training baselines (BGRL [thakoor2022largescale], GraphMAE [hou2022graphmae]) and graph foundation models (OFA [liu2024one], GFT [wang2024gft]). G2PM achieves the best or second-best performance across all datasets, highlighting its superior ability to learn transferable substructure representations, even under significant domain shifts—where message-passing-based methods often struggle.
@@ 6 Related Works
Graph Transformers. Inspired by the success of Transformers in vision and language [devlin2019bert, brown2020language, touvron2023llama, he2022masked, dosovitskiy2021an], several works have extended this architecture to graphs [kreuzer2021rethinking, dwivedi2020generalization, rampasek2022recipe, he2023generalization, chen2023nagphormer, chen2022structure, zhang2024diet, zhang2024mopi]. Typically, graph Transformers treat nodes as tokens and apply self-attention over all pairs. However, the quadratic complexity of node attention limits their scalability, making them impractical for large graphs [wu2023simplifying]. To address this, recent works propose tokenizing graphs via substructures such as random walks or motifs [yanardag2015deep, ivanov2018anonymous], enabling sequence-based modeling [wang2025neural, chen2023nagphormer, he2023generalization, kim2022pure]. For example, GPM [wang2025neural] applies a ViT-style architecture over substructure sequences. Building on this idea, our work explores generative pretraining using a Transformer over substructures, entirely removing message passing.
Generative Pretraining. Generative pretraining has driven major advances in vision and language. These methods predict masked content from visible context, enabling learning from large-scale unlabeled data and powering modern foundation models [touvron2023llama]. This paradigm has extended to other domains (e.g., video, audio, biology), but its impact on graphs remains limited. More recent works apply generative approaches to graphs (e.g., VGAE [kipf2016variational], GraphMAE [hou2022graphmae], GraphMAE2 [hou2023graphmae2], MaskGAE [li2023s], G2PT [chen2025graph]), but focus on reconstructing low-level signals such as nodes or links, and typically rely on message-passing GNNs, limiting their expressiveness and scalability. In contrast, G2PM introduces a fully Transformer-based, message-passing-free framework that models high-level semantic substructures, unlocking better scalability and generalization.
Random Walks on Graphs. Random walks have been widely used to represent substructures, from early unsupervised embeddings [perozzi2014deepwalk, grover2016node2vec, yanardag2015deep, ivanov2018anonymous] to recent deep models [wang2025neural, zhao2022from, jin2022raw, martinkus2023agentbased, tonshoff2023walking, jang2024a, jang2024graph, kim2025revisiting]. These methods have been shown to break key limitations of message passing, improving expressiveness [welke2023expectation, michel2023path], modeling long-range dependencies [martinkus2023agentbased, tonshoff2023walking], and mitigating over-smoothing and over-squashing [wang2024nonconvolutional]. However, these random walk-based methods often focus on global patterns while overlooking localized structures [tonshoff2023walking]. G2PM uses random walks to sample substructures and design the MSM task to automatically balance the contributions between localized and globalized substructures.